OpenStudy (macy):

how do i solve: -5n(n+3)=0 by using the zero-product property?

OpenStudy (anonymous):

-5n=0 n+3=0

OpenStudy (anonymous):

n=0, n=-3

OpenStudy (anonymous):

-5n (n+ 3) =0 -5(2 n) + -15n= 0 -10n + -15n +0 -25n=0 +25 +25 n=25

OpenStudy (macy):

thanks so much!! :):)

OpenStudy (amistre64):

n doesnt equal 25 .... audia was correct clover mistook n*n to mean 2n instead of n^2

OpenStudy (macy):

oh..pellet lol

OpenStudy (macy):

well i still needed the step by step answer anyways.. :/

OpenStudy (macy):

In an exponential model, the y values (?) decrease as the x values increase. a. always b. sometimes c. never

OpenStudy (macy):

please help!

OpenStudy (amistre64):

ill help; but save the vulgarity for a different forum please. we are trying to keep this a family friendly environment....

OpenStudy (amistre64):

sometimes is my best guess

OpenStudy (amistre64):

it depends on the base being used. If the base is between 0 and 1; it increases as x gets small if its greater than 1; it decreases as x gets small

OpenStudy (amistre64):

well, same principle; different x values direction :)

OpenStudy (amistre64):

2^x increases as x gets bigger (1/2)^x decreases as x gets bigger, so sometimes

OpenStudy (macy):

yeah that was my guess too. thanks

OpenStudy (amistre64):

yw :)

OpenStudy (macy):

:)